OpenStudy (anonymous):

Solve 10x^4-27x^2+5=0

OpenStudy (compassionate):

\[x \approx 4472, - 0.4472, 1.5811, 1.5811 \] This is in decimal form.

OpenStudy (anonymous):

is x= +/- sqrt (5/2) correct?

OpenStudy (anonymous):

this is factorable

OpenStudy (compassionate):

x\[x = \sqrt{5}/5\]

OpenStudy (anonymous):

forgot x^2 you'll need to use a u sub though... if you did that, i'm pretty sure you have the right answer

OpenStudy (anonymous):

i got that as well as +/- sqrt 10/2

OpenStudy (anonymous):

\[(2x^2-5)(5x^2-1)=0\] solve i wrote it wrong in the equation editor

OpenStudy (anonymous):

+/ - ^4(sqrt)10 , 3(sqrt)5^2 / 5

OpenStudy (anonymous):

^4(sqrt)10 is the 4th root of 10 where 4 is written above the radicand

OpenStudy (anonymous):

correction: do not place parenthesis around sqrt

OpenStudy (anonymous):

your first answers are right \[+-\sqrt{5}/5\] and \[+-\sqrt{10}/2

OpenStudy (anonymous):

no need for 4th roots i wrote it wrong. i thought you notice must haven ot

OpenStudy (anonymous):

oh sorry